please help me solve:
verify identity:
cosx(tan^2x+1)=secx

Answer by lwsshak3(11628)

verify identity:
cosx(tan^2x+1)=secx
start with left side
cosx[(sin^2x/cos^2x)+1]
=cosx(cos^2x+sin^2x)/cos^2x
=cosx(1)/cos^2x
1/cosx=secx
verified:
left side=right side